1. Mary has a job as waitress at a small restaurant. Tips are pooled in a jar and distributed back to employees at the end of each shift by the manager. These tips are not considered taxable income because they are funded by the patrons rather than the owner of the business. T or F

2. The employer may choose any pay cycle they like. T or F

3. Salaried employees are paid based on the number of hours worked in a pay cycle. T or F

4. Pay cycles are standardized in Canada. T or F

5. Taxable benefits are cash amounts paid to employees in exchange for working. T or F

6. The employment contract may express payment in any method imagined (for example: per hour, per year, per unit, per dollar sold, per telephone call, per page written, etc.).

7. Tips are considered to be a form of earnings. T or F

8. What are the gross weekly earnings for an individual who earns $26, 000 per year and works 35 hours per week?

a. $400

b. Depends on tax rate

c. $1000

d. $800

e. $500

9. An employer with a one-week pay cycle offers Rob $18 per hour and 3 weeks paid vacation in the summer. Calculate his wages for the period if he worked 20 hours.

a. $381.60

b. $360.00

c. $361.20

d. $339.40

e. $410.40

10. The following is not an example of a taxable allowance:

a. Cash paid to an employee to purchase normal clothing to wear on the job.

b. Cash paid to cover living expenses.

c. Cash paid to cover the cost of personal vehicle which may be used for company business.

d. Cash paid to cover travel expenses such as hotels, airfare, taxis, meals, etc.

e. Cash paid to cover entertainment costs not for the benefit of the company.

11. Cell phone service paid for by the employer for the employee's personal use is:

a. Not taxable

b. Taxable with CPP and EI deductions

c. Taxable with CPP deductions but not EI

d. Taxable with EI deductions but not CPP

e. Not taxable but subject to CPP deductions
Answer #1

Answer-

1. False,  All cash tips received by an employee in any calendar month are subject to social security and Medicare taxes and must be reported to the employer, unless the tips received by the employee during a single calendar month while working for the employer total less than $20 . Proper etiquette suggests that patrons should tip between ten to twenty percent of the total bill. ... Tips and other forms of gratuity constitute taxable income and must be reported by those who receive them

2. True, A employer may choose following method of pay cycle , but as per state laws

a, Monthly pay

b. Weekly Pay

c. hourly pay

The Fair Labor Standards Act (FLSA) does not prohibit employers from changing paydays. But the law states that wages must be paid when due, which typically means the next regularly scheduled payday

3. True, a salaried employee may be paid subject to condition enumerated in state labour laws.

A salaried employee is paid based on an annual amount, called a salary. This salary is divided between the pay periods (as determined by the company) for the year and based on a 2080-hour year.

Example: An hourly employee is paid $9.62 an hour. To find this employee's payment amount, the hourly rate is multiplied by the number of hours worked in a pay period.

For calculation purposes, a salaried employee is determined to work 2080 hours a year (52 weeks times 40 hours a week). So, in the examples above, the $9.62 an hour paid to the hourly worker is roughly the same as the $20,000 annual salary paid to the salaried worker, Hence salary may be paid on hourly cirlcle

4. True, Pay cycle are standardized as per employment standards act.

5. False, it include non-cash expense also.

Some of the non-cash nature taxable benefits are-

Tips, boarding, lodging, rent-free or low-rent housing.
travel expenses for personal travel.
personal use of an employer's automobile.
gifts over $500 per year.
use of vacation property owned by the company.
holiday trips.
prizes and awards.

6. True, the employer may refer such payment terms.

Employment contract refers to term of services specified by the employer to employee. Employee can additionally express payment terms on the basis of some terms are as per hour, per year, per unit, per dollar sold, per telephone call, per page written, etc

7. True, Tips are normally be considered as taxable income in the hands of employee as per state laws.

8. Answer is (E) - $ 500

calculation as-

= 26000/(52*35) = $ 14.28571 per hour

weekly rate = 35 * $ 14.28571 = $ 500

9. Answer is (D) $339.40

calculation as :

Regular Weekly earning = 18 * 20 = $ 360

Regular annual Earning = $ 360 * 52 = $ 18720

Weekly off Earning for 3 weeks = $ 360 * 3 = $ 1080

Net Regular annual earning excluding weekly off earning = $ 17640 (18720-1080)

Adjusted weekly earning = $ 17640/52

= $ 339.23

10. Answer is

(e)  Cash paid to cover entertainment costs not for the benefit of the company.

as its not business expense.

11. Answer is (b) Taxable with CPP and EI deductions

Comments- Please give your review, if you liked my answer.